solve 4(x+5)-8x=-2(x-2)

Answer by rkennedy(3)

solve 4(x+5)-8x=-2(x-2)
Distribute
4x + 20 - 8x = -2x + 4
Combine like terms
-4x + 20 = -2x + 4
Subtract 20 from both sides and add -2x to both sides to isolate x on the right
-4x + 20 - 20 = -2x + 4 - 20
-4x = -2x - 16
-4x + (-2x) = -2x - 16 + 2x
-4x - 2x = -16
-2x = -16
divide both sides by -2
-2x/(-2) = -16/-2
x = 8
Check:
4 (8 + 5) - 8(8) = -2 (8 - 2)
32 + 20 - 64 = -16 + 4
52 - 64 = -12
-12 = -12